OpenStudy (anonymous):

Calculate the pH of a Solution which Contain 5 * 10^-8 moles of Hcl in 1 litre of water at 25 C

OpenStudy (anonymous):

@Kryten @Aperogalics

OpenStudy (anonymous):

c= n/V = 5*10^-8 mol/1dm3 = 5*10^-8 M pH = - log [H+] pH = - log(5*10^-8)

OpenStudy (vincent-lyon.fr):

I'm afraid Kryten answered too quickly. The result would be greater than 7, which is not possible for an acidic solution. This is because the H3O+ ions brought by water itself cannot be neglected. Actually, you need to write down water self-ionization reaction and work out the amounts in solution so that they match water's ionic product.

OpenStudy (vincent-lyon.fr):

Answer should be pH = 6.89

OpenStudy (anonymous):

Yup u r correct @Vincent-Lyon.Fr but Hw u Did it ?

OpenStudy (vincent-lyon.fr):

Can you do this ? "write down water self-ionization reaction and work out the amounts in solution" Use notations: \(K_w\) water self-ionization constant \(c_o\) concentration of the strong acid h = \([H_3O^+]\) ω = \([OH^-]\)

OpenStudy (anonymous):

kw = [H+] [0H-]

OpenStudy (vincent-lyon.fr):

OK! Now, write down amounts in solution (1 litre) using the reaction's stoichiometry.

OpenStudy (anonymous):

HCl + H20 ----> H30 + + Cl-

OpenStudy (vincent-lyon.fr):

Ok, This reaction is not an equilibrium and leads to an initial amount of \(c_o\) mole of \(H_3O^+\) in water. Now write down the relevant equation, which is that of the reaction undergone by water.